Table of Contents

2-12-2014

Instrument: Zeiss 1540 (RPI cleanroom)

XC2C32A VQG44 sample was decapped in package using sulfuric acid. Not tested post-decap.

Evaporated amorphous carbon using E&ES system. Target thickness was 20 nm, actual value not measured.

Mounted directly to multi-sample holder using carbon tape. All leads appeared well grounded.

Cut through nitride over three wires, then backfilled holes and deposited 20 um x 20 um Pt pads.

Cleaned for ~5 mins in indirect O2 plasma. C residue remained.

Soldered sample to board and tested. Completely dead, not even JTAG IDCODE response. No current measurements collected.

2-27-2014

Instrument: Zeiss 1540 (RPI cleanroom)

XC2C32A VQG44 sample was decapped in package using sulfuric acid. Tested and confirmed working post-decap. Soldered short stub wires on to aid test.

Evaporated amorphous carbon using E&ES system. Target thickness was 25 nm. Used bare <110> Si wafer as thickness measurement target, measured ~21 nm by ellipsometry.

Mounted to pin on multi-sample holder with carbon tape. Ensured all critical leads were grounded, most others seemed to make good contact but this wasn't tested.

Cut through nitride over one wire, then backfilled hole and deposited 15 x 15 um Pt pad.

Cleaned for ~15 mins in direct O2 plasma, some C residue remained. Repeated for a total of 45 mins at ~100W (standard photoresist strip process). Several pins were likely floating during this process.

Plugged sample into breadboard and tested. Completely dead, not even JTAG IDCODE response. VCCIO+VCCAUX had notably high leakage (20+ mA).

Current hypothesis: Plasma clean is killing chips, try using flood gun on FEI Versa and not coating.

3-27-2014

Instrument:FEI Versa 3D (MRC)

Decapped sample, soldered to board to verify electrical functionality, desoldered.

Mounted on sample holder with carbon tape, put more tape on top of leads to ensure good grounding.

Loaded in chamber, ran setup as per SOP

Initialized both beams, eucentric adjustment. Initial I-beam imaging 10 pA, Focused I-beam at 1 nA on CMP filler. A bit too high, shifted to 300 pA.

Initial Pt dep at 10 pA, used 1 uA for pad.

 
azonenberg/fibnotes.txt · Last modified: 2014/03/27 15:07 by azonenberg
 
Except where otherwise noted, content on this wiki is licensed under the following license: CC Attribution 4.0 International
Recent changes RSS feed Donate Powered by PHP Valid XHTML 1.0 Valid CSS Driven by DokuWiki